On December 31, 2017, Berclair Inc. had 500 million shares of common stock and 23 million shares of 9%, $100 par value cumulative preferred stock issued and outstanding. On March 1, 2018, Berclair purchased 60 million shares of its common stock as treasury stock. Berclair issued a 6% common stock dividend on July 1, 2018. Four million treasury shares were sold on October 1. Net income for the year ended December 31, 2018, was $1,150 million. Also outstanding at December 31 were 90 million incentive stock options granted to key executives on September 13, 2013. The options were exercisable as of September 13, 2017, for 90 million common shares at an exercise price of $78 per share. During 2018, the market price of the common shares averaged $90 per share. The options were exercised on September 1, 2018. Required: Compute Berclair’s basic and diluted earnings per share for the year ended December 31, 2018. (Enter your answers in millions (i.e., 10,000,000 should be entered as 10).)
Answer #1

Weighted Average Oustanding Common Shares

Outstanding Equity Shares on 1st January 2018 = 500 million

Less: Buyback of Outstanding Shares on 1st March 2018 = 60 million = 60 million *10/12 = 50 million

Outstanding after buyback = 500 million - 50 million = 450 million

Stock Dividend on 1st July = 6%

Stock Dividend Multiplier = 450 million * 1.06 = 477 million

New Shares Sold = 4 million on 1st October = 4 million * 3/12 = 1 million

Total Shares Outstanding at the end of the year 2018 = 477 million + 1 million = 478 million

Income for the year 2018 = $1150 million

Basic earning per share = 1150/478 = $ 2.40 per share

Diluted Earning per Share

Potential Shares = Total Options - (Total Oprions * Excercise Price/Average Market Price) = 90 million - (90 million * 78/90) = 12 million

Diluted Outstanding Shares = 478 million + 12 million = 490 million

Diluted Earnings per Share = 1150/490 = $2.35 per share